An aqueous solution containing 3.42 grams per woo grams of water had a freezing point depression of 0.186°c.calculate mass of solute
1
Expert's answer
2019-10-07T07:21:28-0400

"\Delta T=KC_m=K·\frac{m_{salt}}{M_{salt}·m_{H_2O(kg)}}"

"M_{salt}=\frac{K·m_{salt}}{\Delta T·m_{H_2O(kg)}}=\frac{1.86·3.42}{0.186·0.001}=34200g/mol"